If the population of a Fantasy Island is growing at 2% per year and if its population in the year 2006 was 2,200,000.

a. Determine an exponential equation representing its population as a function of time.
b. Determine its population in 2020.
c. Determine what was its population in 2004.

Answer by rothauserc(4718)

The exponential growth or decay formula is
:
a) y(t) = a * e^(kt), y(t) is value at time t, a is the beginning value, k is the rate of growth(k>0) or decay(k<0), t is time, e is euler's number approximately 2.71828
:
b) t = 2020 - 2006 = 14
population(t=14) = 2200000 * (2.71828)^(0.02 * 14) = 2910885.587 = 2,910,886
:
c) t = 2006 - 2004 = 2
population(t=2) = 2200000 * (2.71828)^(-0.02 * 2) = 2113736.76 = 2,113,737
: